OpenStudy (anonymous):

3 cos(theta-60)=2 cos theta

OpenStudy (anonymous):

What to do in this?

OpenStudy (anonymous):

solve plzzzzz

OpenStudy (jhannybean):

hint:\[\cos(a-b) = \cos(a)\cos(b)+\sin(a)\sin(b)\]

OpenStudy (jhannybean):

\[3[cos(\theta -60)]=2\cos(\theta)\]\[3[\cos(\theta)\cos(60) + \sin(\theta)\sin(60)]=2\cos(\theta)\]

OpenStudy (jhannybean):

Started it off for you, go ahead and finish it.

OpenStudy (anonymous):

got it... thaaaaaank you

OpenStudy (jhannybean):

No problem :)